Redan Manufacturing uses 2,500 switch assemblies per week and then reorders another 2,500. The relevant carrying cost per switch assembly per year is $10.00, and the fixed order cost is $ 1,250. Suppose the firm manages its inventory over an one year period. What is the Economic Order Quantity ( EOQ)?
Answer #1
Annual Demand = D = 2500*52 = 130000/year
Ordering Cost = Co = $1250
Holding Cost = Cc = $10
Economic Order Quantity Q* = √(2DCo/Cc) = √(2*130000*1250*/10) = 5700.88